OpenStudy (anonymous):

How do I create a quadratic equation with x = -5 , x = 2

OpenStudy (anonymous):

you need a y =

OpenStudy (lalaly):

y=(x+5)(x-2) =x^2 +3x -10

OpenStudy (anonymous):

There is no y

OpenStudy (anonymous):

That's why I'm confused lol

OpenStudy (lalaly):

a quadratic equation is equal to 0 so x^2 +3x -10 =0

OpenStudy (anonymous):

Sweet! Thank you! =)

OpenStudy (lalaly):

a quadratic function has y